### v2

# PhyGDPO: Physics-Aware Groupwise Direct Preference Optimization for Physically Consistent Text-to-Video Generation

Abstract
Recent advances in text-to-video (T2V) generation have achieved good visual quality, yet synthesizing videos that faithfully follow physical laws remains an open challenge. Existing methods mainly based on graphics or prompt extension struggle to generalize beyond simple simulated environments or learn implicit physical reasoning. The scarcity of training data with rich physics interactions and phenomena is also a problem. In this paper, we first introduce a Physics-Augmented video data construction Pipeline, PhyAugPipe, that leverages a vision–language model (VLM) with chain-of-thought reasoning to collect a large-scale training dataset, PhyVidGen-135K. Then we formulate a principled Physics-aware Groupwise Direct Preference Optimization, PhyGDPO, framework that builds upon the groupwise Plackett–Luce probabilistic model to capture holistic preferences beyond pairwise comparisons. In PhyGDPO, we design a Physics-Guided Rewarding (PGR) scheme that embeds VLM-based physics rewards to steer optimization toward physical consistency. We also propose a LoRA-Switch Reference (LoRA-SR) scheme that eliminates memory-heavy reference duplication for efficient training. Comprehensive experiments show that our method significantly outperforms state-of-the-art open-source methods on the PhyGenBench and VideoPhy2 datasets.
Figure 1: Text-to-video generation on four challenging action categories: (a) gymnastics, (b) soccer, (c) basketball, and (d) glass smashing. When using our post-training method, PhyGDPO, on Wan2.1-T2V-14B [55], the model yields more physically plausible results than OpenAI Sora2 [6] and Google Veo3.1 [18] by generating well-structured human bodies and reasonable physical interactions such as foot kicking the soccer, basketball passing through the hoop, and glass shattering.
1 Introduction
With the development of computing power and the increasing scale of training data, text-to-video (T2V) generation [55, 6, 10] has witnessed significant progress. However, accurately and consistently modeling the physics in the generated video still remains challenging and less explored. Improving the physics reasoning capability of video generation models can make them closer to real-world simulators, which can benefit a wide range of applications such as video gaming [9, 52], autonomous driving [58, 66, 61], robotics [1, 2, 15], film making [48, 57, 62], etc.
To improve the physics modeling, graphics-based methods [29] rely on simulation engines to specify physical parameters of simple scenarios such as perfectly elastic collisions and basic rigid body dynamics. Nonetheless, it is impractical to apply them in real-world scenes as the environments are too complex to parameterize.
Another technical route [69, 64] is based on prompt extension with a large language model (LLM). These methods adopt an LLM to extend the prompts with explicit physics laws and phenomena, and then use the extended prompts as input to simulate the physics by iteratively generating the video or finetuning the model on a small subset. These methods simply follow the LLM-augmented prompts and lack the ability of thinking in physics. Instead, they use the LLM as their surrogate brain and outsource the reasoning process to it. Even so, the prompt following ability of current T2V models is still limited and current LLM’s physics reasoning ability is also weak and often erroneous, which may in turn mislead the T2V model when it follows such guidance to generate videos.
To learn the implicit physics, most current T2V foundation models are trained on massive collections of high-quality text–video data pairs. However, collecting and annotating such data is extremely costly and labor-intensive. Moreover, models obtained through such supervised fine-tuning (SFT) or training still exhibit limited physics reasoning ability. For instance, OpenAI Sora2 [6] and Google DeepMind Veo3.1 [18] often fail in complex human motions or physical phenomena, as shown in Fig. 1, 4, 5, and 6. The key reason is that there are no negative training data to provide contrastive signals discouraging physically inconsistent generations.
The emergence of direct preference optimization [49] (DPO) may provide a potential solution to address the limitations of SFT. However, directly using DPO may encounter three problems: (i) Lack of training data pairs that comprehensively capture physical activities, interactions, and phenomena. (ii) The supervision is mainly based on the Bradley–Terry (BT) probabilistic model, which only compares pairwise data samples. Such binary win/lose feedback lacks the guidance of physics-aware rewarding and struggles to capture inherently holistic global preference signals, such as physical plausibility and motion smoothness. (iii) Vanilla DPO duplicates the training model as a fixed reference, which consumes substantial GPU memory and significantly decreases computational efficiency.
To address these issues, we firstly propose a data construction pipeline, PhyAugPipe, that exploits a vision-language model (VLM) to filter text-video data pairs capturing rich physical interactions and phenomena from a large T2V data pool by parsing the entities and reasoning their actions and outcomes using our designed chain-of-thought (CoT) [60] rule for each data sample. Then we use PhyAugPipe to collect a training dataset, PhyVidGen-135K, containing over 135K text-video data pairs. Secondly, we formulate a principled Physics-aware Groupwise Direct Preference Optimization (PhyGDPO) framework for physically consistent T2V generation. Our PhyGDPO is based on the groupwise Plackett-Luce (PL) probabilistic model, which stimulates the probability distribution over a ranked list of candidate videos and thus enables holistic preference adaptation rather than being constrained to isolated pairwise judgments. Plus, the original DPO training uses the model itself as reward [49] but the model itself can not provide sufficiently good signals and knowledge of physics. Hence, we propose to embed more physics reward during training. Unlike GRPO [51], which suffers from slow speed and unstable optimization, we propose a Physics-Guided Rewarding (PGR) scheme that leverages a physics-aware VLM to guide the DPO data sampling favor more challenging physical actions and allow physics-violating samples to exert stronger influence during preference optimization. To improve the DPO training efficiency and stability, we also propose a LoRA [26]-Switch Reference (LoRA) mechanism that does not need to copy the full model as the reference, which occupies redundant GPU memory. Instead, we attach a trainable LoRA module with a switch to the frozen backbone. Benefit from our dataset and proposed techniques, PhyGDPO significantly improves the physical plausibility of the T2V model, Wan2.1-T2V-14B [55], and yields better results than the two recent best commercial T2V models, OpenAI Sora2 [6] and Google DeepMind Veo3.1 [18] on some challenging action categories, as shown in Fig. 1, 4, 5, and 6.
Our contributions can be summarized as follows:
  • We formulate a principled DPO framework, PhyGDPO, for physically consistent text-to-video generation.
  • We design PGR to guide data sampling and preference optimization favor challenging physical actions. We propose LoRA-SR scheme to reduce GPU memory occupancy for more efficient and stable DPO training.
  • We present PhyAugPipe to construct physics-rich text-video pairs. We use PhyAugPipe to collect a training dataset, PhyVidGen-135K, with over 135K data pairs for studying physically consistent T2V generation.
  • Experiments demonstrate that our PhyGDPO outperforms SOTA methods on the PhyGenBench and VideoPhy2 datasets while yielding higher human preference in the user study of physical realism.
2 Method
Figure 2: Our physics-augmented video data construction pipeline (PhyAugPipe) first adopts a VLM, Qwen-2.5-72B-Instruct [65], following our designed chain-of-thought (CoT) rule in (b) to select text-video data pairs that contain rich physics interactions and phenomena from a large-scale high-quality text-video data pool in (a). Then in (d), we perform action clustering on the filtered data pairs from (c) through the semantics matching via a sentence Transformer [50]. Subsequently, in (e), we adopt a physics-aware VLM, VideoCon-Physics [3], to evaluate the difficulty of different action categories and then sample the text-video pairs accordingly as the wining cases of our training data for preference optimization.
2.1 Physics-Augmented Video Data Construction
Our physics augmented video data construction pipeline (PhyAugPipe) is shown in Fig. 2. We use a VLM, Qwen-2.5-72B-Instruct [65], with our chain-of-thoughts (CoT) in (b) to filter videos with rich physics interactions and phenomena in (c) from a T2V data pool in (a). Then in (d), we cluster the filtered data according to their action categories and sample the data with the rewards of physics-aware VLM, VideoCon-Physics [3], in (e).
Data Filtering with CoT. As shown in Fig. 2 (b), Qwen2.5 with our CoT first parses the elements from the given prompts and video frames, including the physical objects with materials, actions and forces between them. Based on the parsed elements, Qwen2.5 reasons how the entities interact and what results, and rates the physics richness of the data with a score from 0 to 1. Eventually, Qwen2.5 extends the prompt with explicit physics reasoning. Although we explore the implicit physics reasoning ability of T2V models, but our dataset still include the extended prompts to support other research purposes, such as LLM-guided T2V. Please refer to Alg. 1 for the details of our CoT rules. Then in Fig. 2 (c), we threshold the data according to the estimated physics richness.
Algorithm 1 Brief Pipeline of our Chain-of-Thoughts Prompts
1:Original prompt $p$, video frames $V$
2:JSON object containing: original, parse, reason, extended, physics_richness, physics_label
3:
4:Step 1: Element Parsing
5:Extract entities, actions, forces, and outcomes from $(p,V)$ using the VLM.
6:Ensure all agents are included and avoid unsupported or speculative items.
7:Save results into the dictionary parse.
8:
9:Step 2: Vision Checking
10:Compare parse with video frames and the original prompt.
11:Remove hallucinated elements and add missing visible entities or interactions.
12:Update parse accordingly.
13:
14:Step 3: Physics Reasoning
15:Produce a concise explanation describing how the parsed entities interact through physical forces and lead to observed outcomes.
16:Store the explanation as reason.
17:
18:Step 4: Data Scoring
19:Compute physics_richness $\in[0,1]$ based on:
20: – number and interaction of entities;
21: – presence of explicit forces and outcomes;
22: – causal clarity in reason;
23: – penalties from camera motion, stylization, or static-aftermath filters.
24:Set physics_label = 1 if physics_richness $\geq 0.60$, otherwise 0.
25:
26:Step 5: Prompt Extending
27:Extend the original prompt by inserting causal physical details based on reason, without adding new entities, forces, or sensory descriptions.
28:Limit the extension to at most 100 words and store it as extended.
29:
30:return JSON object containing all computed fields.
Action Clustering via Semantics Matching. After data filtering, the retained samples may exhibit distributional imbalance on different actions. To check this issue, we categorize the filtered data into semantically coherent action clusters. Specifically, in Fig. 2 (d), we first compile a list containing $K_{a}$ challenging action categories. Then we feed the original text prompt of each filtered sample and the action list into a sentence Transformer [50] to perform fuzzy semantics matching, thereby determining the action category to which each data sample belongs. We count the number of data samples in each action category to obtain the distribution histogram $\mathcal{H}_{f}$ for the filtered data.
Figure 3: Overview of our PhyGDPO framework. The text prompts without physics reasoning extension in (a) are fed into the T2V model. We propose a LoRA-switch reference scheme in (b) to save the GPU memory and increase the training stability. In (c), PhyGDPO is based on the groupwise Plackett-Luce (PL) probabilistic model and adopts a physics-aware VLM, VideoCon-Physics [3], to reward the DPO training with the real video as the wining sample as it has perfect physics.
Data Sampling with Physics Rewarding. As different action categories vary significantly in physical complexity thus leading to different generation difficulties for the T2V model, we propose to balance the training data accordingly. In particular, we first evaluate how well the T2V model performs on each action category. Based on the performance distribution, we adjust the sampling ratio by allocating more samples to the categories where the model performs poorly. As shown in Fig. 2 (d), we first select the top-$n_{c}$ samples with the highest semantics matching scores within each category as its representatives. Then we employ a physics-aware VLM, VideoCon-Physics [3], to evaluate each video representative. VideoCon-Physics outputs a semantics adherence score and a physics commonsense score $\in[0,1]$ to measure the physics plaucibility of the video. We average the two scores to obtain the overall score for each video. Then we compute the mean score of all representatives within each category to obtain the category score. In this way, we derive the score distribution histogram $\mathcal{S}_{f}$. Then we sample the data according to the difficulty of each action category. More specifically, we define the difficulty of the $k$-th action category as $d_{k}=1-\mathcal{S}_{f}(k)$ and assign a sampling weight following an exponential form as $r_{k}=\text{exp}(\tau d_{k})$, where $\tau$ is a hyperparameter that controls how strongly the sampling favors difficult action categories. Given the total sampling budget $N$, the number of data pairs sampled from the $k$-th action category is formulated as
$\normalsize\mathcal{H}_{r}(k)=\min\bigg(\mathcal{H}_{f}(k),\,N\cdot\frac{r_{k}}{\sum_{j=1}^{K_{a}}r_{j}}\bigg),$ (1)
where $\mathcal{H}_{r}$ denotes the distribution histogram after data sampling with physics rewarding. This sampling strategy allocates more data to challenging actions, encouraging the model to learn more complex physics during training.
2.2 Physics-Aware Groupwise Direct Preference Optimization
Existing DPO algorithms are based on Bradley–Terry (BT) probabilistic model, which compares pairwise data samples, showing limitations in capturing global and holistic preference signals and aligning with human feedback. Besides, previous DPO [54, 36] mainly use the T2V model itself as the reward [49], but the model itself exhibits weak physics reasoning ability as it usually generates physically inconsistent videos. To address these issues, we formulate a Physics-Aware Groupwise Direct Preference Optimization (PhyGDPO), as shown in Fig. 3.
Groupwise Probability. We denote the reward as $r(c,x_{0})$ with the generation $x_{0}$ and condition $c$. Different from the normal DPO, we start from the groupwise Plackett-Luce (PL) probabilistic model, as shown in Fig. 3 (c). We adopt the real video as the winning case $x_{0}^{w}$ because it always follows physical laws and a set of generated videos as the losing cases $\mathcal{G}^{l}(c)=\{x_{0}^{l_{1}},\ldots,x_{0}^{l_{m}}\}$. The preference probability of the PL model is formulated as
$\small p_{\mathrm{PL}}\!\big(x^{w}_{0}\mid\mathcal{G}^{l}(c),c\big)=\frac{\exp\!\big(r(c,x^{w}_{0})\big)}{\exp\!\big(r(c,x^{w}_{0})\big)+\sum_{j=1}^{m}\exp\!\big(r(c,x^{l_{j}}_{0})\big)}.$ (2)
$r(c,x_{0})$ can be parameterized by a neural network $\phi$ [49] and estimated via maximum likelihood training as
$\normalsize\mathcal{L}_{\mathrm{PL}}(\phi)=-\,\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\!\Bigg[\log\frac{\exp\!\big(r_{\phi}(c,x^{w}_{0})\big)}{\exp\!\big(r(c,x^{w}_{0})\big)+\sum_{j=1}^{m}\exp\!\big(r_{\phi}(c,x^{l_{j}}_{0})\big)}\Bigg].$ (3)
Vanilla DPO [36, 54] represents $r(c,x_{0})$ by the T2V model itself as
$\normalsize r(c,x_{0})=\beta\log\frac{p_{\theta}^{*}(x_{0}|c)}{p_{\psi}(x_{0}|c)}+\beta\log Z(c),$ (4)
where $p_{\theta}$ and $p_{\psi}$ denote the conditional probabilities of trained model $\theta$ and reference model $\psi$. $p_{\theta}^{*}$ and $Z(c)$ are the unique global optimal solution and the partition function. We plug Eq. (4) into Eq. (2) and drop the group-constant term $\beta\log Z(c)$ as it does not affect the optimization direction to obtain the groupwise DPO (GDPO) loss as
$\mathcal{L}_{\text{GDPO}}(\theta)$ $=-\,\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\Big[\log\frac{\exp\!\big(\beta f_{\theta}(x^{w}_{0},c)\big)}{\exp\!\big(\beta f_{\theta}(x^{w}_{0},c)\big)+\sum_{j=1}^{m}\exp\!\big(\beta f_{\theta}(x^{l_{j}}_{0},c)\big)}\Big]$ (5)
$=\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\Big[\log\frac{\exp\!\big(\beta f_{\theta}(x^{w}_{0},c)\big)+\sum_{j=1}^{m}\exp\!\big(\beta f_{\theta}(x^{l_{j}}_{0},c)\big)}{\exp\!\big(\beta f_{\theta}(x^{w}_{0},c)\big)}\Big]$
$=\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\Big[\log\!\Big(1+\sum_{j=1}^{m}\exp\!\big(\beta(f_{\theta}(x^{l_{j}}_{0},c)-f_{\theta}(x^{w}_{0},c))\big)\Big)\Big].$
Figure 4: Results of two challenging actions (gymnastics and polo) on VideoPhy2. Our method generates more physically consistent videos, showing coherent, deformation-free gymnastic movements and realistic ball–mallet striking interactions.
where we define the function $f_{\theta}(x_{0},c)$ as the log-likelihood ratio as
$\normalsize\log\frac{p_{\theta}(x_{0}|c)}{p_{\psi}(x_{0}|c)}=\sum_{k=1}^{T}\log\frac{p_{\theta}(x_{t_{k-1}}\mid x_{t_{k}},c)}{p_{\psi}(x_{t_{k-1}}\mid x_{t_{k}},c)}\;\triangleq\;\sum_{k=1}^{T}\Delta_{k}.$ (6)
Here the terminal distribution is inherently represented as the joint transition process over multiple timesteps $t_{k}$ by the definition of diffusion or flow matching models. Subsequently, we reformulate Eq. (5) and estimate its upper bound using the Jensen’s inequality into a single timestep for more efficient training as
$\mathcal{L}_{\text{GDPO}}(\theta)$ $=\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\Big[\log\!\Big(1+\sum_{j=1}^{m}\exp\!\big(\beta T\mathbb{E}_{k}[\Delta_{k}^{l_{j}}-\Delta_{k}^{w}]\big)\Big)\Big]$ (7)
$\leq\mathbb{E}_{c,\mathcal{G}^{l}(c),k}\Big[\log\!\Big(1+\sum_{j=1}^{m}\exp\!\big(\beta T[\Delta_{k}^{l_{j}}-\Delta_{k}^{w}]\big)\Big)\Big].$
Although the upper bound in Eq. (7) allows single-timestep training, it needs to infer the model $2m+2$ times for each group, which requires a long time and large GPU memory. To handle this issue, we exploit an inequality as
$\normalsize 1+\sum_{j=1}^{m}e^{x_{j}}\leq\prod_{j=1}^{m}\big(1+e^{\alpha_{j}x_{j}}\big)^{\gamma_{j}},\penalty 10000\ \penalty 10000\ \penalty 10000\ 0<\alpha_{j}\leq 1,\ \gamma_{j}\geq 1/\alpha_{j}.$ (8)
Please refer to the supplementary for the detailed proof process of this inequality. Subsequently, we can further formulate the upper bound of $\mathcal{L}_{\text{GDPO}}(\theta)$ into a single sample within each group for efficient training as
$\mathcal{L}_{\text{GDPO}}(\theta)$ $\leq\mathbb{E}_{c,\mathcal{G}^{l}(c),k}\Big[\sum_{j=1}^{m}\gamma_{j}\log\Big(1+\exp\big({-\alpha_{j}\beta T[\Delta_{k}^{w}-\Delta_{k}^{l_{j}}]})\big)\Big]$ (9)
$=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-\gamma_{j}\log\sigma\big(-\alpha_{j}\beta T(\Delta_{k}^{w}-\Delta_{k}^{l_{j}})\big)\Big].$
This upper bound allows efficient training with single data pair sampling in a single timestep for each iteration.
Figure 5: Comparison on two challenging random user-input actions (squash and handspring). Our method generates more physically plausible videos, capturin racket–ball interactions in squash and well-coordinated body motion in handspring.
Physics-Guided Rewarding. As aforementioned, normal DPO framework uses the T2V model itself as the reward like Eq. (4), which cannot provide sufficiently good physics knowledge. To embed more physics guidance into our DPO framework, we parameterize $\gamma_{j}$ and $\alpha_{j}$ in Eq. (9) by the normalized semantics adherence score $s^{sa}_{j}$ and physics commonsense score $s^{pc}_{j}$ predicted by a physics-aware VLM, VideoCon-Physics [3], as
$v_{j}$ $=1-\frac{s^{sa}_{j}+s^{pc}_{j}}{2},\penalty 10000\ \penalty 10000\ \gamma_{j}=\frac{1+\lambda\cdot\sigma\!\big(\kappa_{\gamma}(v_{j}-b_{\gamma})\big)}{\alpha_{\text{min}}},$ (10)
$\alpha_{j}$ $=\alpha_{\text{min}}+(1-\alpha_{\text{min}})\cdot\tanh\big(\kappa_{\alpha}(v_{j}-b_{\alpha})\big),$
where $v_{j}$ measures the physics difficulty and mainly modulates $\gamma_{j}$ and $\alpha_{j}$, $\alpha_{\text{min}}$ sets a lower bound to avoid vanishing gradients, $\tanh(\cdot)$ provides a smooth nonlinear mapping that gradually sharpens the preference decision boundary as the gap increases, and the sigmoid function $\sigma(\cdot)$ ensures a bounded and smooth adjustment of $\gamma_{j}$, stabilizing the optimization. $\alpha_{j}$ adaptively adjusts the sharpness of the preference comparison, while $\gamma_{j}$ amplifies the physics-guided reward for samples with lower physical plausibility. They are controlled by the hyperparameters $\lambda$, $\kappa_{\alpha,\gamma}$, $b_{\alpha,\gamma}$, and $\alpha_{\text{min}}$. Our designed physics rewards $(\alpha_{j},\gamma_{j})$ balances the training stability with adaptivity, allowing physics-violating samples to exert stronger influence during optimization.
Flow Matching Probabilistic Modeling. Our method is based on the standard rectified flow matching as
$\normalsize x_{t}=(1-t)\,x_{0}+t\,x_{1},\penalty 10000\ \penalty 10000\ x_{1}\sim\mathcal{N}(0,I),$ (11)
which induces the oracle velocity $u^{*}(x_{t},t\mid c)=x_{1}-x_{0}.$ Discretizing time with step size $h$ (so $t\mapsto t-h$), the backward update is $x_{t-h}=x_{t}-h\,u^{*}(x_{t},t\mid c)$. Since the rectified flow one-step reverse transition has no intermediate noise, we follow a vanishing-noise regularization and approximate $p_{\theta}(x_{t-h}\mid x_{t},c)$ and $p_{\psi}(x_{t-h}\mid x_{t},c)$ as
$p_{\theta}(x_{t-h}\mid x_{t},c)$ $\approx\mathcal{N}\!\big(x_{t}-h\,v_{\theta}(x_{t},t,c),\,\varepsilon I\big),$ (12)
$p_{\psi}(x_{t-h}\mid x_{t},c)$ $\approx\mathcal{N}\!\big(x_{t}-h\,v_{\psi}(x_{t},t,c),\,\varepsilon I\big),\vskip 2.84526pt$
then let $\varepsilon\to 0$ and use the log-ratio between Gaussians with the identity $(x_{t-h}-x_{t})/h=-u^{*}(x_{t},t\mid c)$, we obtain
$\log\frac{p_{\theta}(x_{t-h}\mid x_{t},c)}{p_{\psi}(x_{t-h}\mid x_{t},c)}$ $\approx-\frac{h}{2\varepsilon}\Big(\ell_{\theta}(x_{t},t)-\ell_{\psi}(x_{t},t)\Big),$ (13)
$\ell_{\theta}(x_{t},t)$ $=\|v_{\theta}(x_{t},t,c)-u^{*}(x_{t},t\mid c)\|_{2}^{2},$
$\ell_{\psi}(x_{t},t)$ $=\|v_{\psi}(x_{t},t,c)-u^{*}(x_{t},t\mid c)\|_{2}^{2},$
and reformulate the upper bound of Eq. (9) into the final overall training objective as
$\normalsize\mathcal{L}=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-\gamma_{j}\log\sigma\!\big(-\alpha_{j}\beta T[(\ell_{\theta}^{w}-\ell_{\psi}^{w})-(\ell_{\theta}^{l_{j}}-\ell_{\psi}^{l_{j}})]\big)\Big],\vskip 2.84526pt$ (14)
where any global constant scale factor such as $h/2\varepsilon$ in Eq. (13) can be absorbed into the hyperparameter $\beta$. For simplicity, we denote $\ell_{\theta}^{w}=\ell_{\theta}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\psi}^{w}=\ell_{\psi}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\theta}^{l_{j}}=\ell_{\theta}(x_{t_{k}}^{l_{j}}\mid t_{k})$, and $\ell_{\psi}^{l_{j}}=\ell_{\psi}(x_{t_{k}}^{l_{j}}\mid t_{k})$ in Eq. (14).
LoRA-Switch Reference. Previous DPO methods usually copy the full model and fix it as the reference, which occupies redundant GPU memory and degrades computational efficiency. In addition, this full-copy strategy often leads to unstable and less effective DPO training because this strategy updates the entire set of the backbone weights and may cause the trainable model to quickly and dramatically deviate from the frozen reference model.
Methods Hard Activity Interaction Overall
Vcrafter2 [11] 0.0222 0.1071 0.0943 0.1034
Wan2.1-14B [55] 0.0111 0.1190 0.1572 0.1288
Hunyuan [33] 0.0222 0.1286 0.1572 0.1356
VideoDPO [36] 0.0167 0.1310 0.1572 0.1373
PhyT2V [64] 0.0389 0.1405 0.1698 0.1492
Sora2 [6] 0.0389 0.1429 0.1698 0.1508
Veo3.1 [18] 0.0444 0.1405 0.1887 0.1525
PhyGDPO (Ours) 0.0500 0.1571 0.1761 0.1627
Methods Mechanics Optics Thermal Material Avg
Lavie [59] 0.40 0.44 0.38 0.32 0.36
Wan2.1-14B [55] 0.36 0.53 0.36 0.33 0.40
Open-Sora [73] 0.43 0.50 0.44 0.37 0.44
Pika [47] 0.35 0.56 0.43 0.39 0.44
Vchitect-2.0 [19] 0.41 0.56 0.44 0.37 0.45
PhyT2V [64] 0.45 0.55 0.43 0.53 0.50
VideoDPO [36] 0.48 0.60 0.47 0.58 0.54
PhyGDPO (Ours) 0.55 0.60 0.58 0.47 0.55
Compared Methods Preference of Ours
Vcrafter2 [11] 94.2 %
VideoDPO [36] 89.4 %
PhyT2V [64] 88.5 %
Wan2.1-14B [55] 86.5 %
Hunyuan [33] 82.7 %
Sora2 [6] 67.3 %
Veo3.1 [18] 64.4 %
PhyGDPO (Ours)
Table 1: Quantitative comparison and user preference (%) of our method on VideoPhy2 [4] and PhyGenBench [42] datasets.
Figure 6: Results of two challenging physics phenomena (water refraction and paper combustion) on PhyGenBench [42]. Our method can model the magnifying convex-lens effect and the realistic light refraction caused by the curved water surface, as well as the physically correct ignition and flame propagation on paper, demonstrating stronger physics reasoning ability.
To address this issue, we design a LoRA-Switch Reference (LoRA-SR) scheme. As shown in Fig. 3 (b), we freeze the backbone as the reference model $\psi$ and attach trainable LoRA modules to $\psi$ as the trained model $\theta$. Our LoRA-SR enables the reference and trainable models to share the same heavy backbone while flexibly switching the lightweight LoRA parameters for the training and reference modes. Thus, we can evaluate the trainable and reference terms without storing or loading an extra full model. This significantly increases computational efficiency and prevents the model from drifting too far from the reference, thus stabilizing the training process.
3 Experiment
Dataset. We first use PhyAugPipe to process one million high-quality text-video pairs. After physics data filtering in Fig. 2 (c), we obtain 135K data pairs with sufficient physics interaction and phenomena. Then we perform action clustering and data resampling with physics rewarding to derive 17K data pairs from the 135K filtered data samples. Subsequently, we use the pre-trained T2V model Wan2.1-14B to generate videos for the 17K text prompts with different random seeds and use VideoCon-Physics [3] to score them. We adopt the short unextended prompts from the two datasets, VideoPhy2 [4] and PhyGenBench [42], for evaluation.
[…]
Method Hard Activity Interaction Overall
Baseline 0.0333 0.1405 0.1635 0.1475
+ Chain-of-Thought 0.0389 0.1452 0.1698 0.1525
+ Action Clustering 0.0500 0.1524 0.1698 0.1575
+ Physics Rewarding 0.0500 0.1571 0.1761 0.1627
Method Hard Activity Interaction Overall
Baseline 0.0111 0.1190 0.1572 0.1288
+ LoRA-SR 0.0278 0.1381 0.1635 0.1458
+ Groupwise Model 0.0389 0.1500 0.1698 0.1559
+ Physics Rewarding 0.0500 0.1571 0.1761 0.1627
Method Hard Activity Interaction Overall
Baseline 0.0118 0.1136 0.1447 0.1232
Flow-DPO [35] 0.0296 0.1247 0.1342 0.1283
VideoDPO [36] 0.0278 0.1262 0.1509 0.1305
PhyGDPO (Ours) 0.0444 0.1357 0.1635 0.1407
Method GPU Memory Storage Space Hard Score Overall Score
Baseline - - 0.0118 0.1232
LoRA-SFT 24.7GB 84MB 0.0167 0.1283
w/o LoRA-SR 48.7GB 5.3GB 0.0389 0.1373
with LoRA-SR 25.3GB 84MB 0.0444 0.1407
Table 2: Ablation study on VideoPhy2 [4]. (a) and (b) study the components of PhyAugPipe and PhyGDPO towards better performance. (c) compares VideoDPO with PhyGDPO with the same settings. (d) studies the effect of LoRA-SR scheme.
Figure 7: Visual analysis. (a) When using LoRA-SR, groupwise probabilistic modeling, and physics-guided rewarding, the generated tennis ball motion becomes more physically realistic, better adhering to the laws of buoyancy. (b) Using our PhyGDPO can generate more physically coherent actions with well-structured human body motion than VideoDPO and Flow-DPO. (c) Applying our LoRA-SR can generate more accurate and plausible hand–shovel contact in snow shoveling.
3.1 Comparison with State-of-the-Art Methods
Quantitative Results. We compare PhyGDPO with state-of-the-art (SOTA) methods in Tab. LABEL:tab:videophy2 and LABEL:tab:phygenbench. PhyGDPO outperforms SOTA methods by large margins. (i) Our method surpasses the two recent strongest closed-source models, Sora2 [6] and Veo3 [18] across the tracks of hard actions, activities and sports, and the overall score on VideoPhy2. Especially on the hard actions, PhyGDPO achieves 4.5× higher score than the base model, Wan2.1-14B, and is 29% and 13% higher than Sora2 and Veo3. (ii) Compared to the SOTA DPO algorithm for video generation (VideoDPO) and SOTA method for physically consistent video generatiion (PhyT2V), our method surpasses VideoDPO and PhyT2V and by 200% and 29% on the hard action score of VideoPhy2. Our method is also 22%/15% and 35%/23% higher than PhyT2V/VideoDPO on the mechanics and thermal tracks of PhyGenBench.
User Study. Since the auto-evaluation tools of VideoPhy2 and PhyGenBench are based on VLMs, their assessment of physical plausibility may be inaccurate. Therefore, we conduct a user study with 104 participants. Each participant views the videos generated by our PhyGDPO and a random competing method, along with the randomly selected text prompt from VideoPhy2 and PhyGenBench. The participants are asked to pick up the video that better follows the physics laws. Tab. LABEL:tab:user_study reports the user preference (%) of our method over competing entries. PhyGDPO is consistently preferred by the human evaluators than all SOTA methods.
Qualitative Results. The visual comparisons on challenging physical actions or phenomena including gymnastics, soccer, basketball, glass smashing, polo, squash, handspring, refraction, and combustionare shown in Fig. 1, 4, 5, and 6. Our method generates videos with more realistic physical dynamics, deformation-free body motion, and accurate object interactions, effectively capturing phenomena such as force transfer, material deformation, light refraction, and flame propagation. Compared to existing open- and closed-source models, PhyGDPO shows superior generalization from human activities to complex physical events. Please refer to the project page for video results.
3.2 Ablation Study
Ablation of PhyAugPipe. We conduct a break-down ablation to study the effect of each component of PhyAugPipe in Tab. LABEL:tab:phyaugpipe. The baseline is directly training PhyGDPO on the randomly selected text-video data. When we keep the number of selected training data the same and apply the data filtering with CoT, action clustering with semantics matching, and data sampling with physics-guided rewarding, all track scores are improved. Especially the score on the hard actions gains by over 50%. These results suggest the effectiveness of our data construction techniques.
[…]
PhyGDPO vs. SOTA DPO. For fair comparison with two SOTA DPO methods: Flow-DPO [35] and VideoDPO, we use them to finetune Wan2.1-1.3B with the same data and settings. As reported in Tab. LABEL:tab:dpo_compare, PhyGDPO surpasses Flow-DPO and VideoDPO across all tracks by large margins, especially on the hard action track, where it yields 50% improvement. We conduct a visual comparison in Fig. 7 (b) on the challenging weightlifting action. Flow-DPO and VideoDPO generate distorted or unstable body poses. In contrast, PhyGDPO generates coherent body motion with stable shapes and balanced force dynamics, highlighting its advantage in modeling complex physical actions.
[…]
4 Related Work
4.1 Text-to-Video Generation Models
T2V generation [6, 68, 24, 74, 5, 23, 5, 17, 27, 25, 71, 59] has witnessed significant progress. Many existing works follow DiT [46] to adopt a Transformer [53] to predict the noise in diffusion [41, 34, 39, 40, 20, 48, 8, 22, 32] or estimate the velocity field in flow matching [55, 27, 17, 21, 7, 28, 72, 56, 70, 31, 13]. Although high visual quality is achieved, accurately modeling the underlying physics-related effects remains challenging. Two recent works, DiffPhy [69] and PhyT2V [64], adopt an LLM to extend the text prompt with explicit physics laws and phenomena, and then iteratively generates the video or finetunes a T2V model with extended prompts. Yet, these prompt extension-based methods are easily misled by the mistakes of LLM and struggle to learn the implicit physics. Different from these works, we aim to explore the physics reasoning potential of the T2V model itself.
4.2 Direct Preference Optimization
Direct Preference Optimization (DPO) [49] is proposed to align LLMs with human preferences and has been adopted in image [54, 43, 16, 30] and video [36, 12, 35, 63, 67] generation. For example, DiffusionDPO [54] finetunes a text-to-image diffusion model with human-annotated preference image pairs. VideoDPO [36] adapts DiffusionDPO into T2V generation. However, these DPO frameworks mainly improve the visual quality and aesthetics. In addition, these DPO methods suffer from a severe low-efficiency problem because they need to copy a full model as the reference and they also lack sufficient guidance of physics, which makes the preference signal less reliable and limits the improvement on physical consistency. Our work aims to cope with these research problems.
5 Conclusion
In this paper, we focus on studying a challenging problem, physically consistent T2V generation, without using LLM for prompt extension in inference because our goal is to explore the implicit physics reasoning ability of the video generation model. To this end, we first propose a data construction method, PhyAugPipe, to collect a training dataset PhyVidGen-135K containing over 135K text-video data pairs with rich physics interaction and phenomena. Subsequently, we formulate a principled groupwise DPO framework, PhyGDPO, with two technical designs PGR and LoRA-SR. PhyGDPO efficiently post-trains Wan2.1-T2V-14B model on PhyVidGen-135K to boost the physics plausibility in video generation. Comprehenseiv experiments show that our method quantitatively and qualitatively outperforms SOTA algorithms and achieves higher human preference in the user study.